OpenStudy (anonymous):

What is the value for x in the proportion x over x plus ten equals seven over twelve?

OpenStudy (anonymous):

so its like x/(x+12) =7/12. => 12x =(x+12)*7 => 12x-7x=84. 5x =84. x= 84/5=16.8 :)

OpenStudy (anonymous):

thank you :)